OpenStudy (anonymous):

Monica booked a trip to Disneyland for herself and her family.The total cost of the trip was 5800 pounds plus VAT at 20%. She paid 1300 pounds of the total cost when she booked her holiday. She paid the rest of the total cost in 6 equal months.Work out the amount of each monthly payment,

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute the total cost, first

OpenStudy (michele_laino):

total cost is 5800 + VAT, namely: \[\Large 5800 + \left( {5800 \cdot \frac{{20}}{{100}}} \right) = ...?\]

OpenStudy (texaschic101):

is VAT taxes ??

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

gotcha....I understand now

OpenStudy (texaschic101):

sorry...continue explaining Michele

OpenStudy (anonymous):

1160 @Michele_Laino

OpenStudy (michele_laino):

if you want please continue @texaschic101 :)

OpenStudy (anonymous):

is that correct @Michele_Laino

OpenStudy (anonymous):

please proceed @Michele_Laino

OpenStudy (michele_laino):

yes!, nevertheless you have compute this: 5800+1960=...?

OpenStudy (texaschic101):

total cost would be 5800 + 20% of 5800 5800 + .2(5800) = total cost

OpenStudy (michele_laino):

correct! @texaschic101

OpenStudy (anonymous):

hmm is the answer to this part 1160?

OpenStudy (texaschic101):

not quite.. 5800 + 1160 = total cost

OpenStudy (anonymous):

ok im confused

OpenStudy (texaschic101):

turn your percent to a decimal 20% = 0.2 5800 + 0.2(5800) = total cost 5800 + 1160 = total cost 6960 = total cost understand ?

OpenStudy (anonymous):

yes thanks next please @texaschic101

OpenStudy (texaschic101):

so she paid 1300 6960 - 1300 = ?? (you can do the math)

OpenStudy (anonymous):

5660

OpenStudy (texaschic101):

ok....correct....so we have 5660 she then paid 6 equal monthly installments....so we divide by 6 5660/6 = each monthly installment

OpenStudy (anonymous):

943.333333

OpenStudy (texaschic101):

which rounds to what ?

OpenStudy (anonymous):

940

OpenStudy (anonymous):

or just 943

OpenStudy (texaschic101):

I would probably round it like this.. 943.33333 = 943.33

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

give medal to Michele....I shouldn't have taken over her job at explaining it to you

OpenStudy (anonymous):

ok if you insist but thank you so much @texaschic101

OpenStudy (texaschic101):

no problem :)